OpenStudy (anonymous):

e^2x − 5^ex − 36 = 0

OpenStudy (anonymous):

do u guys know how to solve this?

zepdrix (zepdrix):

This is the problem...? \[\Large\rm e^{2x}-5^{ex}-36=0\]

zepdrix (zepdrix):

Did you mean to write,\[\Large\rm e^{2x}-5e^x-36=0\]

OpenStudy (anonymous):

yes that is the problem

zepdrix (zepdrix):

Rules of exponents let's us rewrite the first term,\[\Large\rm \left(e^{x}\right)^2-5e^x-36=0\]Then we can make a substitution to help us get through this,\[\Large\rm u=e^x\]Giving us,\[\Large\rm u^2-5u-36=0\]

zepdrix (zepdrix):

It looks like it's factor-able from there. Understand how to proceed?

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

its not factorable because 5 doesn't go into 36?

zepdrix (zepdrix):

So we need factors of -36, they have to add to -5. -6*6=-36 but -6+6 = 0. Hmm those didn't work. 9*-4=-36 but 9+-4=5. Hmm that didn't give us -5 but maybe we're on the right track with 9 and 4?

OpenStudy (anonymous):

i didn't get it because it needs to be a decimal number

zepdrix (zepdrix):

That's because we're not done yet. This is only half of the problem. You need to be able to factor in order to get past this part though.